OpenStudy (anonymous):

Simplify the expression. A. 0 B. 1 C. 10 D. 100

OpenStudy (anonymous):

No expression so it must be 0 right?!

OpenStudy (sidsiddhartha):

lol

OpenStudy (anonymous):

10 to the power of 0 sorry my computer being stupid

OpenStudy (anonymous):

When dealing with exponents, when ever you have anything to the power of 0 it equals 1.

OpenStudy (anonymous):

\[2^0 = 1,~~~235235^0 = 1\]

OpenStudy (anonymous):

so my problem would be 1?

OpenStudy (anonymous):

I should write \[(235235)^0 = 1\]

OpenStudy (anonymous):

And yes

OpenStudy (anonymous):

ok thanks